Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
amagazzu - 15:39 Tuesday 12 January 2021 (50421) Print this report
Operator Report - Morning shift

The shift was dedicated to standard Maintenance. Here a list of the activities carried out:

- Standard Vacuum Maintenance ( VAC Team );
- Cleaning inside the experimental Area ( Ciardelli with external firms );
- DAQ Maintenance Activities ( see logbook entry #50417 );
- Increase of Fiber Laser Power ( see #50420 );
- Auto-restart implementation for ALS 2W optical amplifiers ( see #50418 );
- Recovery of SR Suspensions loops after yesterday night earthquake ( carried out by Ruggi from remote );
- DET cryotrap refill operation updated ( see #50419 ).

During the morning, it was also possible to relock the cavities. At 7:45 UTC they were both locked again, but the power was lower than usual. This behavior was caused by the drift control not engaged due to the flag ASC_ENABLE being at 0 ( see entry #50377 ) since yesterday at 15:30 UTC. After restoring its value via Vpm the power in the cavities returned to its nominal values.

Air Conditioning
Under request of M. Mantovani at 11:05 UTC I restored the PyHVAC Servos for INJ and DET.

Comments to this report:
mantovani - 17:08 Tuesday 12 January 2021 (50424) Print this report

Now the HVAC for the labs works properly,see attached plot

Images attached to this comment
Search Help
×

Warning

×